OpenStudy (anonymous):

what is the fractional form of 7.2???? my answer is 72/1000 but idk if fractional form is the same thing as asking convert 7.2 percent into 72/1000 ????

ganeshie8 (ganeshie8):

I think fractional form just means to write \(7.2\) as ratio of two integers : \(\dfrac{a}{b}\)

OpenStudy (anonymous):

7 and 2/10

ganeshie8 (ganeshie8):

\[7.2 = \dfrac{72}{10} = \dfrac{36}{5}\]

OpenStudy (anonymous):

so it would be 36/5 or 72/10????

ganeshie8 (ganeshie8):

72/10 is same as 36/5 but we say 36/5 is simplified form. so use 36/5

OpenStudy (anonymous):

either one is technically correct

OpenStudy (anonymous):

oh oak ill try it thank u i think i get it now

OpenStudy (anonymous):

i mean ok

OpenStudy (jdoe0001):

\(\bf 7.\underline{2}\implies \cfrac{7.2}{1\underline{0}}\) you put at the bottom, "as many zeros as there are decimals" notice, 1 decimal, thus 1 0

OpenStudy (jdoe0001):

well.. and you'd lose the dot also =) \(\bf 7.\underline{2}\implies \cfrac{72}{1\underline{0}}\)

OpenStudy (freckles):

or you can look at it as if you are multiplying numerator and denominator by 10 \[7.2 =\frac{7.2}{1} \cdot \frac{10}{10}\]

OpenStudy (anonymous):

oh ok multiplying numerators helped me understand thanks:)

OpenStudy (freckles):

so if you have \[6.04=\frac{6.04}{1} \cdot \frac{100}{100}=\frac{604}{100} \text{ where this fraction can be reduced } \]